OpenStudy (anonymous):

A biologist is comparing the growth of a population of flies per week to the number of flies an iguana will consume per week. He has devised an equation to solve for which day (x) the iguana would be able to eat the entire population. The equation is 3x = 2x + 1. However, he has observed that the iguana cannot eat more than eight flies in one week. Explain to the biologist how he can solve this on a graph using a system of equations. Identify any possible constraints to the situation.

OpenStudy (solomonzelman):

please fix your equation.

OpenStudy (solomonzelman):

So far I can only say, `iguana cannot eat more than eight flies in one week` is going to give an equation, \(\normalsize\color{blue}{ f(g)≤8x}\) and \(\normalsize\color{blue}{ f(1)≤1 \times 8~~~~~⇒~~~~~f(1)≤8}\) \(\normalsize\color{blue}{ f(2)≤2 \times 8~~~~~⇒~~~~~f(2)≤16}\) \(\normalsize\color{blue}{ f(3)≤3 \times 8~~~~~⇒~~~~~f(2)≤24}\) and on.... this is just showing that, not more than 8 flies per week.